The mass of the great pyramid is 557t greater than that of the leaning tower of pisa. Stone Henge has a mass of 2695t, which is 95t less than the leaning tower of pisa, there once was a greater pyramid which had the mass twice that of the great pyramid. what was the mass of the greater pyramid?
Thankyou

Answer by wuwei96815(245)

Word problems are usually linguistic problems because once you figure out what is being asked the math is normally routine. Let's try to rewrite the word problem as mathematical expressions.

x=mass of Pisa in tons
x+557=great pyramid
x-95=2695 (mass of Stonehenge)
2(x+557)=greater pyramid

Let's solve for x by using the equation x-95=2695:
x-95=2695
x=2695+95=2790

Now let's use our value for x to substitute and solve our problem.
2(x+557)=greater pyramid
2(2790+557)=greater pyramid
6694 tons = greater pyramid

Did we get it right?